Resolve a force of 100 N into two components which lie on opposite sides of the force and each of which makes an angle of 300 with the force.
1
Expert's answer
2020-10-16T10:59:30-0400
"F_{1y}=F_{2y}=0.5F=0.5(100)=50\ N\\F_{1x}=-F_{2x}=\frac{F_{1y}}{\tan{30}}=\frac{50}{\tan{30}}=86.6\ N"